Suzanne borrowed a certain amount of money at an annual simple interest rate of 4.3%, if she returned $5891 after 245 weeks, how much interest did she pay? Round your answer to the nearest dollar.
Answer #1

Answer

245 weeks = 245/52 = 4.71 years

Simple Interest rate =4.3%

P + (P x 4.3% x 4.71) = 5891

Solving P=4899

Interest = $5891 -$4899 =$992.43

Rounded off $ 992